# Fair Play in the Newsroom: Actor-Based Filtering Gender Discrimination in Text Corpora

### v1

@@ Abstract
Large language models are increasingly shaping digital communication, yet their outputs often reflect structural gender imbalances that originate from their training data. This paper presents an extended actor-level pipeline for detecting and mitigating gender discrimination in large-scale text corpora. Building on prior work in discourse-aware fairness analysis, we introduce new actor-level metrics that capture asymmetries in sentiment, syntactic agency, and quotation styles. The pipeline supports both diagnostic corpus analysis and exclusion-based balancing, enabling the construction of fairer corpora. We apply our approach to the taz2024full corpus of German newspaper articles from 1980 to 2024, demonstrating substantial improvements in gender balance across multiple linguistic dimensions. Our results show that while surface-level asymmetries can be mitigated through filtering and rebalancing, subtler forms of bias persist, particularly in sentiment and framing. We release the tools and reports to support further research in discourse-based fairness auditing and equitable corpus construction.
@@ 1 Introduction
Large Language Models (LLM) are increasingly integrated into everyday digital services, from search engines and translation tools to conversational agents and content recommendation systems. Despite their impressive capabilities, LLMs are known to perpetuate and even amplify harmful societal biases, including gender stereotypes, occupational hierarchies, and asymmetric visibility of social groups Torrielli (2025); Armstrong et al. (2024); Siddique et al. (2024); Jeoung et al. (2023). These patterns are not inherent to the models’ architecture itself (Vaswani et al., 2017) but are learned from the data on which they are trained. As training data plays a central role in shaping the representational landscape of LLMs’ outputs, critical analysis of large-scale text corpora becomes a key step towards mitigating downstream harms. Without an understanding of the social structures embedded in textual data, efforts to ensure fairness and accountability in LLMs risk addressing symptoms rather than causes.
Previous work by Urchs et al. (2025) introduced a linguistically grounded pipeline to detect gender discrimination in German newspaper texts through actor-level discourse analysis. Their approach identifies named actors and analyses how they are represented in terms of visibility and associated language, drawing on concepts such as nomination and predication from critical discourse studies.
Building on this foundation, we present an extended pipeline that enables both fine-grained fairness auditing and corpus-level discrimination reduction. Our contribution is fourfold:
@@
    We introduce novel actor-level discrimination markers, including syntactic roles, quote attribution, and sentiment bias.
@@
    We enhance interpretability through structured, human-readable reports that enable qualitative and diachronic analysis.
@@
    We propose a method for generating gender-balanced corpus recommendations by excluding disproportionately discriminatory texts.
@@
    We release the entire pipeline as open-source software to support transparency, reproducibility, and collaborative development.
This paper offers tools and insights for creating fairer NLP systems by revealing how social groups are represented in text. We combine discourse-informed fairness analysis with scalable corpus processing to enable actor-level discrimination detection and targeted corpus balancing.
@@ 2 Related Work and Conceptual Background
Detecting gender discrimination in text requires an interdisciplinary foundation that integrates perspectives from linguistics, gender studies, and computer science. In this section, we first outline our conceptual understanding of gender and linguistic discrimination. We then review existing approaches to computational discrimination detection. Finally, we introduce the actor-level discrimination detection pipeline by Urchs et al. (2025), which forms the foundation of our work.
@@ 2.1 Gender and Linguistic Discrimination
Gender is treated as a socially constructed and co-constructed identity, rather than a fixed biological or grammatical category. While linguistic gender relates to grammatical rules and biological sex is often seen as binary and immutable, we work with the concept of social gender, which can be fluid, non-binary, and shaped through interaction and recognition. This perspective foregrounds the performative and contextual nature of gender, acknowledging its entanglement with social norms, symbolic power, and habitualised practices West and Zimmerman (1987); Ainsworth (2015); Konishi (1993); Kramer (2020); Devinney et al. (2022). However, due to methodological constraints and the structure of the German language, the empirical analysis remains limited to the binary spectrum.
Discrimination, in contrast to bias or fairness, is understood here as a social effect: It is the observable outcome of differential treatment based on protected attributes such as gender. Drawing on Reisigl (2017) and the functional model by Graumann and Wintermantel (2007), we conceptualise linguistic discrimination as an act or process that disadvantages (or favours) individuals through patterns in language. This includes acts of naming (nomination) and describing (predication) actors in ways that reflect or reinforce social hierarchies. Such acts need not be intentional, but they may also arise from the habitual reproduction of dominant norms in public discourse. This framework stands in contrast to many machine learning–based approaches, where bias often refers to statistical imbalances in data or model performance, and fairness is operationalised via metrics such as demographic parity or equal opportunity Blodgett et al. (2020); Caton and Haas (2024). While these frameworks are useful for quantifying group disparities, they tend to reduce complex social identities to binary categories, thereby risking the oversight of contextual, structural, and discursive forms of inequality.
@@ 2.2 Computational Discrimination Detection
Discrimination in text is not limited to overtly offensive statements but often emerges from subtle and structural patterns of language use. Computational discrimination analysis aims to make such patterns visible by identifying systematic disparities in how individuals or groups are represented in text. In computer science, discrimination is often formalised via the concept of fairness and operationalised using statistical metrics such as demographic parity, equalised odds, or individual fairness Mehrabi et al. (2021). These metrics are well-suited for classification tasks and support scalability and reproducibility. However, they also reduce social categories to fixed, binary attributes and abstract away from contextual and structural forms of inequality Blodgett et al. (2020).
When applied to text, this paradigm produces approaches that focus primarily on hate speech detection, sentiment disparity, or stereotyping, often based on keyword lists or supervised classification. Although these methods yield valuable insights, they rarely address how discrimination is embedded in discourse or how it is distributed across different individuals within a text. Instead, most existing systems operate at a text-level granularity, assigning a global label such as “discriminatory” or “non-discriminatory” to entire documents.
In contrast, Urchs et al. (2025) focus on discrimination at the actor level. By identifying individual actors and examining how they are named (nomination) and described (predication), their approach reveals asymmetries in representation within the same text. This enables a more granular and discourse-aware analysis of discrimination in language.
@@ 2.3 Actor-Level Discrimination Detection Pipeline
Our pipeline builds upon prior work by Urchs et al. (2024, 2025). The first paper introduces actor-based fairness analysis in isolated texts using a modular pipeline that combines information extraction with linguistic discourse analysis. Specifically, it detects gender discrimination on the actor level by identifying the nomination and the predication of individual actors. The pipeline extracts actors via named entity recognition (NER), resolves pronouns through coreference resolution, and stores all actor references, including names, titles, and generic forms, in a structured knowledge base. For each actor, all sentences containing them are extracted and analysed for sentiment, gender-coded language, and linguistic framing. Discrimination metrics are computed per actor and summarised into a discrimination report. The report includes a range of linguistic and structural metrics designed to capture different facets of gendered representation:
  • Actor counts: Number of distinct male-, female-, and undefined-coded actors per text.
  • Mention counts: Total number of pronoun or name-based references per gender group.
@@
Their second paper (Urchs et al., 2025) scales this analysis to a large newspaper corpus (taz2024full) with over 1.8 million articles published between 1980 and 2024. It adapts the actor-level pipeline for German, replacing the English sentiment model with a BERT-based classifier trained on German news, and introduces additional markers for gender-neutral language and generic masculine usage. The analysis aggregates actor-based metrics by year, enabling the study of historical shifts in gender representation and framing.
@@
However, the approach remains purely descriptive and diagnostic: No mechanism is implemented for correcting or filtering the corpus based on the findings. Our work extends the work from both papers substantially in depth and scope. We expand the actor-level analysis with new discrimination detection metrics and integrate the pipeline into a two-stage exclusion framework.
@@ 3 The Extended Actor-Centred Pipeline
We extend the original actor-centred pipeline, improving both the granularity of the analysis and its ability to support fairer corpus construction.
Building on insights from systemic functional linguistics Halliday (2004) and critical discourse analysis Reisigl (2017), the pipeline now incorporates syntactic role annotation to capture how subject and object positions contribute to gendered representations. In discourse, actors in subject roles typically perform actions, while those in object positions are acted upon. Tracking these roles across pronoun groups helps reveal patterns of agency and passivity that are central to linguistic discrimination. To deepen the understanding of the gendered representation, we track whether actors are referred to by their name or just by their pronouns. We refined the quote attribution by distinguishing direct from indirect speech using punctuation and reporting verbs. Furthermore, zooming in on how the actors’ utterances are presented gives insights into how active they are in the text. Direct quotes indicate a higher degree of activity than indirect ones. We enrich the framing context by extending pointwise mutual information (PMI) calculations to include not only adjectives but also verbs and nouns. This expansion highlights thematic associations and role-specific language tied to gendered actors. Human-readable reports are redesigned to improve structure, interpretability, and accessibility. All results are presented in a structured plain-text report, with separate sections for summary statistics, syntactic roles, sentiment, and lexical framing. Breakdowns on a pronoun-group basis and PMI tables are aligned and consistently formatted, making the output interpretable even for non-technical readers.
Finally, the pipeline supports a two-step corpus filtering mechanism: Articles exhibiting strong internal gender asymmetries are flagged using multiple indicators, and a subsequent balancing step adjusts overall gender ratios in the dataset. This process yields a curated corpus suitable for training language models with reduced gender discrimination.
The full pipeline code is available on our git repository https://github.com/Ognatai/corpus_balancing.
@@ 4 Pipeline Application: Discrimination Analysis and Corpus Balancing
This section outlines how we implemented the extended actor-level pipeline in two consecutive stages.
4.1 Stage 1: Discrimination Analysis Across the Full Corpus
For the initial discrimination analysis, we build on the pipeline proposed by Urchs et al. (2025) (cf. §2.3). We extend this approach by incorporating a set of additional linguistic and structural metrics that allow for a more fine-grained assessment of gendered representation and framing. These metrics are extracted at the actor level and aggregated per document and year:
  • Mentions: Captures how often an actor is referred to by name (e.g., Angela Merkel) or pronoun (e.g., she, he), enabling analysis of individuation, visibility, and referential strategies.
  • Syntactic roles: Counts how often the actor appears as grammatical subject (nsubj) or object (obj), providing a proxy for discursive agency and passivity.
  • Quotation style: Differentiates between direct and indirect speech attributions, reflecting variation in narrative presence and framing control.
  • Top 10 PMI terms: Lists the nouns and verbs most strongly associated with the actor based on Pointwise Mutual Information (PMI), offering insight into typical roles, actions, and semantic contexts.
The full taz2024full corpus is analysed both at the yearly level and aggregated in total. We also introduce an improved yearly reporting format that organises both existing and newly introduced metrics in a more accessible structure. An example of this yearly report (for 2023) can be found in Appendix A.
4.2 Stage 2: Article Filtering and Corpus Balancing
To reach the endeavour of constructing a more balanced corpus and substantially reducing the impact of gender-discriminating articles, we introduce a multi-stage filtering pipeline. In contrast to stage 1, which analyses the corpus on a yearly basis, this stage operates on the entire dataset to enable global exclusion and balancing decisions.
Step 1: Full Corpus Analysis
The full corpus is processed using the pipeline from Stage 1 (cf. §2.3). Instead of aggregating metrics per year, actor-level knowledge bases and article-level metadata are saved for each article. These intermediate files are reused in the subsequent steps to minimise redundant executions of pipeline steps. After processing, a histogram is generated to visualise the distribution of gender ratios across articles (cf. Figure 5). Each article is assigned two values: the percentage of mentions and the percentage of actors associated with she/her pronouns. The left subplot displays the mention-weighted distribution, and the right subplot shows the actor-weighted distribution. Both histograms use percentage values ranging from 0 (only he/him references) to 100 (only she/her references), with each bar representing the proportion of articles falling into that range.
Step 2: Text-Level Exclusion
An interactive filtering interface allows the user to define how many of the four heuristic exclusion criteria (see list below) must be met to classify a text as discriminatory. The default threshold is two out of four. The criteria are:
  • Sentiment disparity: A large gap in average sentiment scores between female- and male-coded actors.
  • Grammatical role asymmetry: A strong difference in subject-to-object ratios between pronoun groups.
  • Quote attribution imbalance: A pronounced disparity in direct versus indirect quotes.
  • Naming versus pronoun imbalance: Disproportionate usage of named versus pronominal references between genders.
Each criterion is calculated using Laplace-smoothed ratios to avoid instability possibly originating from low-frequency counts. The user may customise the threshold for each criterion at runtime. By default, a text exhibits significant imbalance if the sentiment gap exceeds 0.3, or if the difference in subject/object roles, quote attribution, or naming/pronoun usage exceeds 0.5. These thresholds correspond to cases where one gender group is at least twice as prominent as the other in a specific framing dimension. A lower threshold is used for sentiment because sentiment values tend to cluster around neutral and vary within a narrower range; even small differences may indicate meaningful affective bias. Overall, the chosen defaults strike a balance between interpretability and selectivity, capturing strongly biased texts while preserving as much of the corpus as possible. Articles that trigger the user-defined number of flags are excluded from the corpus. Exclusion decisions are logged in a structured exclusion file and visualised through an updated gender ratio distribution plot (cf. Figure 6).
Step 3: Corpus-Level Balancing
After text-level filtering, we apply corpus-level balancing based on the relative contribution of each article to the overall gender ratio. The user sets an equilibrium range for actor and mention ratios between female- and male-coded references. The default range is $[0.75,1.25]$, allowing for up to 25% deviation in either direction. Articles are iteratively excluded based on their contribution to the global imbalance until both ratios fall within the specified interval. A visualisation of the resulting gender ratio distribution is generated to illustrate the effects of balancing (cf. Figure 7.) All excluded articles are recorded in a structured exclusion file for transparency and reproducibility.
Step 4: Corpus Reconstruction
All excluded article IDs from both filtering stages are consolidated and used to construct a new balanced corpus. Articles are removed directly from the original JSON files, and the revised dataset is saved to disk.

### v2

@@ Abstract
Language corpora are the foundation of most natural language processing research, yet they often reproduce structural inequalities. One such inequality is gender discrimination in how actors are represented, which can distort analyses and perpetuate discriminatory outcomes. This paper introduces a user-centric, actor-level pipeline for detecting and mitigating gender discrimination in large-scale text corpora. By combining discourse-aware analysis with metrics for sentiment, syntactic agency, and quotation styles, our method enables both fine-grained auditing and exclusion-based balancing. Applied to the taz2024full corpus of German newspaper articles (1980–2024), the pipeline yields a more gender-balanced dataset while preserving core dynamics of the source material. Our findings show that structural asymmetries can be reduced through systematic filtering, though subtler biases in sentiment and framing remain. We release the tools and reports to support further research in discourse-based fairness auditing and equitable corpus construction.
@@ 1 Introduction
Large-scale text corpora are central to natural language processing and related fields, yet they often reproduce societal inequalities. Wikipedia reflects gender imbalances in coverage Wagner et al. (2021), job advertisements use gendered wording that reinforces hierarchies Gaucher et al. (2011), and the film industry promotes stereotypes Kagan et al. (2020). Such examples show how corpora encode and normalise discrimination in persistent ways. Detecting these patterns is essential, but given the scale of modern datasets, manual inspection is infeasible. Automatic methods are needed to reveal structural inequalities at scale, and crucially, detection must be paired with curation: once problematic material is identified, corpora should be rebalanced to provide more reliable input data for NLP applications and more trustworthy resources for research.
Urchs et al. (2025) introduced a linguistically grounded pipeline to detect gender discrimination in German newspapers through actor-level discourse analysis, examining how named actors are represented via nomination and predication. Building on this work, we extend the pipeline to enable both fine-grained fairness auditing and corpus-level discrimination reduction. Our contributions are:
@@
    Novel actor-level discrimination markers, including syntactic roles, quote attribution, and sentiment bias.
@@
    Structured, human-readable reports that support qualitative and diachronic analysis.
@@
    A method for generating gender-balanced corpora by excluding disproportionately discriminatory texts.
@@
    An open-source release of the pipeline to ensure transparency, reproducibility, and collaboration.
This paper offers tools and insights for creating fairer corpora by revealing how social groups are represented in text. We combine discourse-informed analysis with scalable processing to enable actor-level discrimination detection and targeted corpus balancing.
@@ 2 Related Work and Conceptual Background
Detecting gender discrimination in text requires an interdisciplinary foundation that integrates perspectives from linguistics, gender studies, and computer science.
@@ 2.1 Gender and Linguistic Discrimination
Gender is treated here as a socially constructed identity rather than a fixed biological or grammatical category. While linguistic gender follows grammatical rules Kramer (2020), and NLP research often reduces gender to binary labels Devinney et al. (2022), we work with the broader notion of social gender, which is fluid, contextual, and shaped through interaction West and Zimmerman (1987). Our empirical analysis is restricted to binary categories because the corpus lacks sufficient non-binary representation, but the approach can be adapted to encompass more inclusive forms of gender representation.
Discrimination, in contrast to bias or fairness, is understood here as a social effect: the observable outcome of differential treatment based on protected attributes such as gender. Following Reisigl (2017), we view social discrimination as a process that disadvantages individuals through recurring patterns in language. This perspective differs from many machine learning approaches, where bias is framed as statistical imbalance and fairness as compliance with formal metrics such as demographic parity or equal opportunity Blodgett et al. (2020); Caton and Haas (2024). While effective for measuring distributional disparities, these frameworks largely ignore semantic and discursive aspects of language, where subtle forms of discrimination are often embedded.
@@ 2.2 Computational Discrimination Detection
In computational research, discrimination is usually formalised through fairness metrics such as demographic parity, equalised odds, or individual fairness Mehrabi et al. (2021). While these approaches are scalable and reproducible, they treat social categories as fixed attributes and largely abstract away from semantics and discourse Blodgett et al. (2020). Applied to text, this has produced methods for hate speech detection, sentiment disparity, or stereotyping, typically relying on keyword lists or supervised classifiers. Such methods yield valuable insights but operate mainly at the document level, labelling texts as“discriminatory” or “non-discriminatory” and overlooking how unequal treatment is distributed within discourse.
In contrast, Urchs et al. (2025) proposed an actor-level approach that identifies individuals and analyses how they are represented through nomination and predication. By shifting the focus from entire texts to the representation of actors within them, this perspective reveals structural asymmetries that remain invisible to classical bias-detection methods.
@@ 2.3 Actor-Level Discrimination Detection Pipeline
Our pipeline builds on prior work by Urchs et al. (2024, 2025). The first paper introduces actor-based fairness analysis in isolated English texts using a modular pipeline that combines information extraction with discourse analysis. It detects gender discrimination at the actor level by identifying nomination and predication, extracting actors via named entity recognition (NER), resolving pronouns through coreference, and storing references (names, titles, generic forms) in a structured knowledge base. For each actor (per text), all sentences in which they are mentioned are analysed for sentiment, gender-coded language, and framing. The resulting discrimination report provides per-text metrics such as:
  • Actor counts: Number of distinct male-, female-, non-binary- and undefined-coded actors per text.
  • Mention counts: Total number of pronoun or name-based references per gender group. [^1] The difference between actor counts and mention counts can be illustrated with a simple example: a text with one male actor mentioned ten times differs from a text with ten female actors each mentioned once. Both cases result in ten actor references, but the distribution of visibility is fundamentally different.
@@
The second paper scales this analysis to the taz2024full corpus (1.8M German newspaper articles, 1980–2024). It adapts the pipeline for German, replaces the sentiment model with a BERT-based classifier trained on German news, and adds markers for gender-neutral language and generic masculine usage. Actor-level metrics are aggregated by year, enabling longitudinal analysis of representation and framing. Additional features include:
@@
This approach, however, remains purely descriptive. Our work extends it substantially: we introduce new actor-level discrimination metrics and integrate a two-stage exclusion framework to move from diagnosis to corpus correction.
@@ 3 The Extended Actor-Centred Pipeline
We extend the actor-level pipeline introduced by Urchs et al. (2024) and scaled in Urchs et al. (2025) to improve both analytical granularity and corpus curation. Unlike classical bias detection methods, which rely on document-level labels or aggregate statistics, our approach captures discrimination at the level of individual actors, making structural inequalities within texts visible.
Building on systemic functional linguistics Halliday (2004) and critical discourse analysis Reisigl (2017), the pipeline incorporates metrics targeting key dimensions of discursive inequality:
  • Syntactic roles: Distinguishing subject and object positions provides a proxy for agency. Actors in subject roles are framed as active agents, while object roles position them as passive. Tracking this distribution across gender groups highlights structural asymmetries in agency Halliday (2004).
  • Naming vs. pronoun reference: Whether actors are referred to by name or reduced to pronouns affects their individuation and visibility. Persistent differences between genders can signal unequal treatment in how actors are foregrounded Bendel Larcher (2015).
  • Quotation style: Direct quotations attribute voice and authority, while indirect quotations background speakers. Measuring the ratio of direct to indirect speech shows how discursive authority is distributed Bendel Larcher (2015).
  • Sentiment: The evaluative framing of actors, captured via sentiment analysis, indicates whether certain groups are systematically associated with more negative language.
  • Pointwise Mutual Information (PMI): By extracting strongly associated adjectives, verbs, and nouns, we reveal the thematic and lexical contexts in which actors are embedded, surfacing stereotypical associations.
These metrics go beyond frequency counts to capture framing, which classical fairness metrics (e.g., demographic parity) overlook. Actor-level analysis adds value over methods such as hate speech classification or keyword-based stereotype detection by revealing who is made visible, who is granted agency or voice, and how evaluations differ across gender groups. Insights that word- or document-level approaches cannot provide.
The pipeline outputs structured reports that combine these metrics with summary statistics, enabling both qualitative and quantitative inspection. It also supports a two-stage user-centred filtering mechanism: (1) flagging articles with strong internal asymmetries, and (2) rebalancing overall gender ratios. This ensures that the resulting corpus is not only analysed but also curated to reduce discriminatory patterns. The full pipeline code, documentation, and yearly reports are available at https://anonymous.4open.science/r/corpus_balancing-4531.
@@ 4 Pipeline Application: Discrimination Analysis and Corpus Balancing
We apply the extended actor-centred pipeline in two stages: first for diagnostic analysis, then for corrective balancing. Detection alone is insufficient: if left uncorrected, strong asymmetries risk skewing corpus statistics and reinforcing discriminatory patterns in downstream applications. Our pipeline, therefore, combines analysis with systematic filtering and balancing.
4.1 Stage 1: Discrimination Analysis Across the Corpus
In the first stage, the pipeline processes all articles and computes the actor-level metrics described in Section 2.3 and Section 3. Results are aggregated per article and year to enable both fine-grained inspection and diachronic analysis. Yearly reports combine the full set of metrics in a structured, interpretable format, supporting both quantitative tracking of trends and qualitative exploration of framing practices (see Appendix A for an example). Actors are only tract per text, not in the whole corpus.
4.2 Stage 2: Multi-Stage Filtering and Corpus Balancing
The pipeline first produces a histogram showing, for all articles, the proportion of actors coded with she/her pronouns and the proportion of their mentions, ranging from 0% (only he/him) to 100% (only she/her). This initial view allows users to inspect the distribution of gender ratios before any intervention and to set thresholds for four asymmetry indicators introduced in Section 2.3: sentiment gap, subject/object ratio, quote imbalance, and naming imbalance (named versus pronoun mentions).
Each indicator is computed per article for the two groups and compared as a ratio difference with +1 Laplace smoothing to stabilise small counts. Concretely, for group $g\in\{\textit{she},\textit{he}\}$ we define
$\textit{subject/object}(g)=\frac{\textit{subjects}_{g}+1}{\textit{objects}_{g}+1}$ (1)
$\textit{direct/indirect}(g)=\frac{\textit{direct}_{g}+1}{\textit{indirect}_{g}+1}$ (2)
$\textit{named/pronoun}(g)=\frac{\textit{named}_{g}+1}{\textit{pronoun}_{g}+1}$ (3)
and sentiment(g) is the article level average polarity for mentions of group $g$. An article is flagged on a given indicator if the absolute difference between the two group-specific values exceeds a user-chosen threshold. Users also specify the minimum number of indicators that must be triggered simultaneously for an article to be excluded.
Thresholds are chosen with two principles in mind. First, sentiment operates in a narrow numeric range around neutrality, so even moderate absolute differences are meaningful for evaluative framing. Second, the structural ratios subject/object, direct/indirect, and named/pronoun exhibit higher natural variability across topics and genres, therefore stricter cut offs help avoid false positives from incidental fluctuations. Intuitively, a large ratio difference marks a sustained structural tilt, for example a pattern where one group appears predominantly as grammatical subjects relative to objects, is quoted directly rather than paraphrased, or is referred to by name rather than by pronoun, compared with the other group. This configuration enables flexible yet principled flagging, and Section 5 reports the concrete threshold values used in this study together with their empirical motivation.
After text-level exclusion, a second histogram is generated to show the updated distribution of gender ratios across articles. At this stage, the user decides on an equilibrium range for corpus-level balancing by specifying lower and upper bounds (e.g., how much more men can appear than women, and vice versa).
Corpus-level balancing then iteratively excludes articles that contribute most to the remaining imbalance until actor- and mention-based ratios fall within the chosen range. A final histogram visualises the adjusted distribution and documents the effect of the balancing step.
Finally, all excluded article IDs are consolidated, and a revised balanced corpus is created. It is saved in the same format as the original dataset, but as a new version, ensuring compatibility while providing a fairer foundation for downstream use.